• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name brads

    Log entry time 08:44:13 on March 16, 2012

    Entry number 366658

    This entry is a followup to: 366606

    keyword=Trigger rate anomolies

    Once the target is polarized, we are going to do a BCM calibration. This will allow a DAQ expert to verify that the triggers are working as needed for the start of production running.

    See Halog 366606 and Halog 366649. I find it odd that the T1 is so close to T2, T3:T4 is a more typical ratio. It is also notable that the ratio of R:L singles triggers (T1:T3) is 100:1. Is this all due to the target field (seems unlikely to me, but...)

    A copy of this log entry has been emailed to: vasulk, camsonne, melissac, brads